# A physics-compliant diagonal representation for wireless channels parametrized by beyond-diagonal reconfigurable intelligent surfaces

### v1

@@ IV-B3 Results
Based on the same data set, we ran our end-to-end channel estimation algorithm twice. Two examples of estimated parameter sets are plotted in Fig. 4A and compared to the known ground truth values. Except for the $\mathcal{RT}$ block of $\mathbf{S}^{\mathrm{casc}}$ (which is a scalar in our case study), clear differences to the ground truth as well as between the two estimates are apparent, as expected due to the inevitable ambiguities. Nonetheless, the estimated set of parameters allows us to accurately predict $h(\mathbf{b})$ for arbitrary unseen $\mathbf{b}$, as evidenced by the comparison between predicted and ground-truth values of $h(\mathbf{b})$ for arbitrary unseen $\mathbf{b}$ in Fig. 4B. The inevitable ambiguities are thus not problematic for our goal of identifying an accurate mapping $\mathbf{b}\rightarrow h(\mathbf{b})$. We have hence successfully determined a “purely physics-based digital twin” for our BD-RIS parametrized channel without having explicitly identified $\mathbf{S}^{\mathrm{RE}}$ or $\mathbf{S}^{\mathrm{LC}}$. In fact, our method does not even require knowledge of $N_{\mathrm{S}}$.
Figure 4: (A) Amplitude and phase of the experimentally grounded ground truth (blue) and two separate estimates (red and green) of the parameters within the diagonal representation of the BD-RIS-parametrized channel: $r_{2}-r_{1}$, $r_{1}$, the $\mathcal{RT}$ block of of $\mathbf{S}^{\mathrm{casc}}$ (which is just a scalar here), the $\mathcal{AC}$ block of of $\mathbf{S}^{\mathrm{casc}}$ (which is vectorized for ease of display here), and the $\mathcal{CC}$ block of of $\mathbf{S}^{\mathrm{casc}}$ (which is vectorized for ease of display here). (B) Amplitude and phase of ground truth (blue) and predicted (red) end-to-end wireless channel $h$ for 100 random unseen configurations of the reconfigurable load network.
@@ IV-C RSSI Optimization
The RSSI enhancement enabled by the considered BD-RIS is moderate in the considered example due to the small scale of the RIS which comprises only six elements. However, this example is sufficient to demonstrate the key point of our paper, namely that the we can successfully optimize our BD-RIS based on the end-to-end channel estimate from the previous section. Both channel estimation and optimization pivotally rely on the diagonal representation of the BD-RIS-parametrized channel.
@@
Figure 5: Probability density function of the RSSI for the considered SISO link, as well as predicted and ground-truth RSSI for the configuration $\mathbf{b}_{\mathrm{opt}}$ obtained with Algorithm 1 based on the end-to-end channel estimate from Sec. IV.
V Relation to PhysFad
Thus far, we have only discussed circuit-theory-based physics-compliant modeling of BD-RIS parametrized channels. For the D-RIS case, an alternative (but ultimately equivalent) physics-compliant approach based on a coupled-dipole formalism exists. It was introduced as PhysFad [18] and constituted the basis of the first experimental validation of a physics-compliant channel model for arbitrarily complex D-RIS-parametrized rich-scattering channels as well as various frugal physics-compliant end-to-end channel estimation protocols [12]. The basic approach of the coupled-dipole formalism consists in assigning a dipole to each primary wireless entity (antennas and tunable lumped elements). The dipoles are characterized by their polarizability which is tunable in the case of the tunable lumped elements. The dipoles are coupled to each other via background Green’s functions. An interaction matrix is set up, composed of the sum of a diagonal matrix containing the inverse polarizabilities and a symmetric matrix containing the background Green’s functions. The top left block of the inverse of this interaction matrix yields $\tilde{\mathbf{S}}$. Whereas circuit theory treats systems as black boxes, the coupled-dipole approach describes the physical wave-matter interactions inside the systems, enabling insights that are not accessible via circuit theory such as the method for performing conjugate beamforming without pilot exchange purely based on user motion that was proposed and experimentally demonstrated in Ref. [49].
To date, it remained unclear how exactly the PhysFad formalism could be applied to BD-RIS-parametrized wireless channels. Given the presented physics-compliant diagonal representation of BD-RIS-parametrized channels, the compatibility with the PhysFad framework is now clear: The dipoles are associated with the antennas and tunable lumped elements, and the background Green’s functions account for the potentially highly complex scattering in the cascade of the radio environment and the load circuit.
VI Conclusion
To summarize, we have introduced a new diagonal representation of BD-RIS-parametrized channels that is fully physics-compliant. Our key insight relates to the realization that the BD-RIS reconfigurable impedance network, irrespective of its detailed architecture, is itself a multi-port network with a subset of auxiliary ports terminated by individual tunable loads. This implies that a BD-RIS-parametrized channel results from a chain cascade of radio environment (RE), load circuit (LC) and individual tunable loads (IL). Whereas conventionally one considers the auxiliary ports of RE associated with RIS elements to be terminated by a load that is the cascade of LC and IL (and hence characterized by a generally beyond-diagonal matrix), we alternatively consider the cascade of RE and LC whose auxiliary ports associated with tunable load elements are terminated by IL (which is always characterized by a diagonal matrix). A key consequence of this diagonal representation is that we can directly apply D-RIS algorithms for channel estimation and optimization to BD-RIS-parametrized channels. We demonstrated this in the first experimentally grounded case study of BD-RIS with an example of physics-compliant end-to-end channel estimation and RSSI maximization. Importantly, we only sought a representation of the characteristics of the cascade of RE and LC, without ever requiring its breakdown into RE and LC, and without even requiring a removal of ambiguities. Hence, operationally, there is in principle no need to explicitly know $\mathbf{S}^{\mathrm{RE}}$ and $\mathbf{S}^{\mathrm{LC}}$; nonetheless, for pedagogical reasons, we also demonstrated how to obtain $\mathbf{S}^{\mathrm{LC}}$ analytically for canonical ideal reconfigurable impedance networks, and we discussed how to obtain it numerically or experimentally for practical realizations. Along the way, we have furthermore rigorously derived the relation between physics-compliant and simplified cascaded models of BD-RIS-parametrized arbitrarily complex (i.e., potentially rich-scattering) channels, as well as how the latter can be represented by coupled-dipole-based physics-compliant models like PhysFad.
@@ VI Conclusion
In this Appendix, we illustrate the procedure for analytically determining $\mathbf{S}^{\mathrm{LC}}$ for an ideal T network whose impedances have been replaced by auxiliary ports, as sketched in Fig. 2B. The characteristic impedances of the transmission lines attached to the ports are assumed to be $Z_{0}=50\ \Omega$. We follow the procedure outlined in Pozar’s book [50]: “$S_{ij}$ is found by driving port $j$ with an incident wave of voltage $V^{+}_{j}$ and measuring the reflected wave amplitude $V^{-}_{i}$ coming out of port $i$. The incident waves on all ports except the $j$th port are set to zero, which means that all ports should be terminated in matched loads to avoid reflections.” We also detail various sanity checks to confirm that our results are correct.
-A Step-by-step procedure
@@
    $S_{ii}^{\mathrm{LC,T}}=\frac{V_{i}^{-}}{V_{i}^{+}}=\frac{V_{i}-V_{i}^{+}}{V_{i}^{+}}=\frac{V_{i}-(V_{g}/2)}{V_{g}/2}.$ (16)
@@
    $S_{ji}^{\mathrm{LC,T}}=\frac{V_{j}^{-}}{V_{i}^{+}}=\frac{V_{j}}{V_{g}/2}$ (17)
@@
Figure 6: (A) Ideal T network with auxiliary ports replacing the three impedances. Polarities are defined for each of the five ports. (B) Setup to determine column $i=1$ of $\mathbf{S}^{\mathrm{LC}}$. All ports except for the $i$th port are terminated with matched loads ($Z_{0}$). A matched source (voltage generator in series with a matched load, drawn in light blue) is connected to the $i$th port. The voltage drops across all ports are monitored with the voltmeters drawn in green.
-B Sanity checks
For the considered problem, $\mathbf{S}^{\mathrm{LC,T}}$ must satisfy two basic properties:
@@
    $\left(\mathbf{S}^{\mathrm{LC,T}}\right)^{T}=\mathbf{S}^{\mathrm{LC,T}}.$ (18)
@@
    $\left(\mathbf{S}^{\mathrm{LC,T}}\right)^{\dagger}\mathbf{S}^{\mathrm{LC,T}}=\mathbf{I}_{5}.$ (19)
Verifying that these two expectations are met constitutes the first two sanity checks, which are easily passed by the form of $\mathbf{S}^{\mathrm{LC}}$ defined in Eq. (12).
@@
This ideal T network corresponds to terminating ports 3, 4 and 5 of our 5-port network (sketched in Fig. 2B) with load impedances $Z_{1}$, $Z_{2}$ and $Z_{3}$. Our third sanity check hence consists in verifying that based on the form of $\mathbf{S}^{\mathrm{LC,T}}$ defined in Eq. (12) we recover the well-established $\mathbf{Z}^{\mathrm{T}}$ from Eq. (20) for random complex-valued choices of the values of $Z_{1}$, $Z_{2}$ and $Z_{3}$.
@@
$\mathbf{S}^{\mathrm{T}}=\mathbf{S}^{\mathrm{LC,T}}_{\bar{\mathcal{S}}\bar{\mathcal{S}}}+\mathbf{S}^{\mathrm{LC,T}}_{\bar{\mathcal{S}}\mathcal{C}}\left(\left(\mathrm{diag}(\mathbf{r})\right)^{-1}-\mathbf{S}^{\mathrm{LC,T}}_{\mathcal{CC}}\right)^{-1}\mathbf{S}^{\mathrm{LC,T}}_{\mathcal{C}\bar{\mathcal{S}}},$ (21)
@@
The form of $\mathbf{S}^{\mathrm{LC,T}}$ defined in Eq. (12) passes all three sanity checks.
@@ Acknowledgment
The author acknowledges stimulating discussions with M. Nerini, H. Prod’homme, and A. Shaham.

### v2

@@ IV-B3 Results
Based on the same data set, we ran our end-to-end channel estimation algorithm twice. Two examples of estimated parameter sets are plotted in Fig. 4A in red and green, and compared to the known ground-truth values plotted in blue. Except for the $\mathcal{RT}$ block of $\mathbf{S}^{\mathrm{K}}$ (which is a scalar in our case study), clear differences to the ground truth as well as between the two estimates are apparent, as expected due to the inevitable ambiguities. Nonetheless, the estimated set of parameters allows us to accurately predict $h(\mathbf{b})$ for arbitrary unseen $\mathbf{b}$, as evidenced by the comparison between predicted and ground-truth values of $h(\mathbf{b})$ for arbitrary unseen $\mathbf{b}$ in Fig. 4B. The inevitable ambiguities are thus not problematic for our goal of identifying an accurate mapping $\mathbf{b}\rightarrow h(\mathbf{b})$. We have hence successfully determined a “purely physics-based digital twin” for our BD-RIS parametrized channel without having explicitly identified $\mathbf{S}^{\mathrm{RE}}$ or $\mathbf{S}^{\mathrm{SLC}}$. In fact, our method does not even require knowledge of $N_{\mathrm{S}}$.
Figure 4: End-to-end BD-RIS-parametrized channel estimation in the rich-scattering environment from Fig. 3A using a physics-compliant D-RIS algorithm. The parameters to be estimated are $r_{\mathrm{A}}$ and $r_{\mathrm{B}}$, as well as the blocks $\mathcal{RT}$, $\mathcal{AC}$ and $\mathcal{CC}$ of $\mathbf{S}^{\mathrm{K}}$. Given these parameters, for any given $\mathbf{b}$ we can evaluate $\mathbf{S}^{\mathrm{IL}}$ using Eq. (20) and then $h$ using Eq. (10). (A) Amplitude and phase of the experimentally grounded ground truth (blue) and two separate estimates (red and green) of the estimated parameters. For ease of visualization, we display the matrix-valued blocks $\mathcal{AC}$ and $\mathcal{CC}$ of $\mathbf{S}^{\mathrm{K}}$ in vectorized form. (B) Amplitude and phase of ground-truth (blue) and predicted (red) end-to-end wireless SISO channel $h$ for 100 random unseen configurations of the BD-RIS in the rich-scattering environment from Fig. 3A.
@@ IV-C RSSI Optimization
The RSSI enhancement enabled by the considered BD-RIS is moderate in the considered example due to the small scale of the RIS which comprises only six elements. However, this example is sufficient to demonstrate the key point of our paper, namely that the we can successfully optimize our BD-RIS based on the end-to-end channel estimate from the previous section. Both channel estimation and optimization pivotally rely on the diagonal representation of the BD-RIS-parametrized channel in Eq. (10).
@@
Figure 5: Probability density function of the RSSI for the considered SISO link, as well as predicted and ground-truth RSSI for the configuration $\mathbf{b}_{\mathrm{opt}}$ obtained with Algorithm 1 based on the end-to-end channel estimate from Sec. IV-B.
V Conclusion
To summarize, we have introduced a new diagonal representation of BD-RIS-parametrized channels that is fully physics-compliant. Our key insight is based on the realization that the BD-RIS reconfigurable impedance network can be separated into a static part and a set of individual tunable loads. This implies that a BD-RIS-parametrized channel results from a chain cascade of radio environment (RE), static parts of the load circuit (SLC) and individual tunable loads (IL). By considering the cascade K of the static and non-diagonal systems RE and SLC, which is terminated by the tunable diagonal system IL, we achieve a representation that is directly analogous to physics-compliant models for D-RIS-parametrized channels. Therefore, our diagonal representation enables us to directly apply physics-compliant D-RIS algorithms to scenarios with BD-RIS-parametrized channels. We have evidenced this fact by using a physics-compliant D-RIS algorithm to perform end-to-end channel estimation and optimization in an experimentally grounded case study of a BD-RIS-parametrized rich-scattering radio environment.
Along the way, we have formalized the operational equivalence of physics-compliant models for D-RIS based on multi-port network theory and the coupled-dipole formalism (PhysFad). This equivalence, combined with our diagonal representation for BD-RIS within multi-port network theory, has allowed us to fill a research gap on how to describe BD-RIS-parametrized channels within PhysFad. Moreover, for pedagogical purposes, we have discussed in detail how one can determine the scattering parameters of SLC in various scenarios arising in theoretical, numerical and experimental studies. However, we have emphasized that explicitly knowing the characteristics of SLC is operationally not necessary. Indeed, our diagonal representation for BD-RIS-parametrized channels is the most compact physics-compliant representation and its deployment does not require one to explicitly separate K into RE and SLC.
Altogether, we expect our physics-compliant diagonal representation for BD-RIS-parametrized channels to enable a paradigm shift in how practitioners in wireless communications and signal processing implement system-level optimizations of BD-RIS-parametrized radio environments. Specifically, our findings obviate the need for developing system-level optimization algorithms dedicated to BD-RIS, because existing physics-compliant algorithms already developed for D-RIS can be applied directly.
Our work also has a take-away message regarding the performance comparison between D-RIS and BD-RIS. Currently, such comparisons always fix the number of RIS elements. Our work points to the importance of comparative studies comparing for a fixed number of tunable lumped elements rather than a fixed number of RIS elements. Indeed, our work highlights that the primary entities in a BD-RIS-parametrized environment are the antennas and the tunable lumped elements in the load circuit but not the RIS elements themselves; indeed, knowledge of the number of RIS elements is not even required when using our physics-compliant diagonal representation of BD-RIS-parametrized channels, as seen in our case study.
Appendix A Procedure to determine $\mathbf{S}^{\mathrm{SLC,T}}$ analytically
@@
In this Appendix, we illustrate the procedure for analytically determining $\mathbf{S}^{\mathrm{SLC}}$ for an ideal T network whose impedances have been replaced by auxiliary ports, as sketched in Fig. 2B. The characteristic impedances of the transmission lines attached to the ports are assumed to be $Z_{0}=50\ \Omega$. We follow the procedure outlined in Pozar’s book [53]: “$S_{ij}$ is found by driving port $j$ with an incident wave of voltage $V^{+}_{j}$ and measuring the reflected wave amplitude $V^{-}_{i}$ coming out of port $i$. The incident waves on all ports except the $j$th port are set to zero, which means that all ports should be terminated in matched loads to avoid reflections.” We also detail various sanity checks to confirm that our results are correct.
A-A Step-by-step procedure
@@
    $S_{ii}^{\mathrm{SLC,T}}=\frac{V_{i}^{-}}{V_{i}^{+}}=\frac{V_{i}-V_{i}^{+}}{V_{i}^{+}}=\frac{V_{i}-(V_{g}/2)}{V_{g}/2}.$ (23)
@@
    $S_{ji}^{\mathrm{SLC,T}}=\frac{V_{j}^{-}}{V_{i}^{+}}=\frac{V_{j}}{V_{g}/2}$ (24)
@@
Figure 6: (A) Ideal T network with auxiliary ports replacing the three impedances. Polarities are defined for each of the five ports. (B) Setup to determine column $i=1$ of $\mathbf{S}^{\mathrm{SLC}}$. All ports except for the $i$th port are terminated with matched loads ($Z_{0}$). A matched source (voltage generator in series with a matched load, drawn in light blue) is connected to the $i$th port. The voltage drops across all ports are monitored with the voltmeters drawn in green.
A-B Sanity checks
For the considered problem, $\mathbf{S}^{\mathrm{SLC,T}}$ must satisfy two basic properties:
@@ A-B Sanity checks
    $\left(\mathbf{S}^{\mathrm{SLC,T}}\right)^{T}=\mathbf{S}^{\mathrm{SLC,T}}.$ (25)
@@
    $\left(\mathbf{S}^{\mathrm{SLC,T}}\right)^{\dagger}\mathbf{S}^{\mathrm{SLC,T}}=\mathbf{I}_{5}.$ (26)
Verifying that these two expectations are met constitutes the first two sanity checks, which are easily passed by the form of $\mathbf{S}^{\mathrm{SLC}}$ defined in Eq. (18).
@@
This ideal T network corresponds to terminating ports 3, 4 and 5 of our 5-port network (sketched in Fig. 2B) with load impedances $Z_{1}$, $Z_{2}$ and $Z_{3}$. Our third sanity check hence consists in verifying that based on the form of $\mathbf{S}^{\mathrm{SLC,T}}$ defined in Eq. (18) we recover the well-established $\mathbf{Z}^{\mathrm{T}}$ from Eq. (27) for random complex-valued choices of the values of $Z_{1}$, $Z_{2}$ and $Z_{3}$.
@@
$\mathbf{S}^{\mathrm{T}}=\mathbf{S}^{\mathrm{SLC,T}}_{\bar{\mathcal{S}}\bar{\mathcal{S}}}+\mathbf{S}^{\mathrm{SLC,T}}_{\bar{\mathcal{S}}\mathcal{C}}\left(\left(\mathrm{diag}(\mathbf{r})\right)^{-1}-\mathbf{S}^{\mathrm{SLC,T}}_{\mathcal{CC}}\right)^{-1}\mathbf{S}^{\mathrm{SLC,T}}_{\mathcal{C}\bar{\mathcal{S}}},$ (28)
@@
The form of $\mathbf{S}^{\mathrm{SLC,T}}$ defined in Eq. (18) passes all three sanity checks.
Appendix B Derivation of Eq. (16)
For notational ease, we use $\mathbf{\Phi}=\mathrm{diag}(\mathbf{c})$, $\mathbf{\Sigma}=\left[\mathbf{W}^{0}_{\mathcal{SS}}\right]^{-1}$, $\mathbf{\Omega}(\mathbf{c})=\mathbf{W}^{-1}(\mathbf{c})$ and $\mathbf{\Omega}^{0}=\mathbf{W}^{-1}(\mathbf{c}=\mathbf{0})$. The block-matrix inversion lemma directly yields
$\mathbf{\Omega}^{0}_{\mathcal{AA}}=\left(\mathbf{W}^{0}_{\mathcal{AA}}-\mathbf{W}^{0}_{\mathcal{AS}}\mathbf{\Sigma}\mathbf{W}^{0}_{\mathcal{SA}}\right)^{-1},$ (29a)
$\mathbf{\Omega}^{0}_{\mathcal{AS}}=-\mathbf{\Omega}^{0}_{\mathcal{AA}}\mathbf{W}^{0}_{\mathcal{AS}}\mathbf{\Sigma}=\left[\mathbf{\Omega}^{0}_{\mathcal{AS}}\right]^{T},$ (29b)
$\mathbf{\Omega}^{0}_{\mathcal{SS}}=\mathbf{\Sigma}+\mathbf{\Sigma}\mathbf{W}^{0}_{\mathcal{SA}}\mathbf{\Omega}^{0}_{\mathcal{AA}}\mathbf{W}^{0}_{\mathcal{AS}}\mathbf{\Sigma}.$ (29c)
Moreover, the Woodbury matrix identity yields the following auxiliary result:
$\left(\mathbf{W}^{0}_{\mathcal{SS}}-\mathbf{\Phi}\right)^{-1}=\mathbf{\Sigma}-\mathbf{\Sigma}\left(-\mathbf{\Phi}+\mathbf{\Sigma}\right)^{-1}\mathbf{\Sigma}.$ (30)
According to the block-matrix inversion lemma,
$\left[\mathbf{\Omega}(\mathbf{c})\right]_{\mathcal{AA}}=\left[\mathbf{W}^{0}_{\mathcal{AA}}-\mathbf{W}^{0}_{\mathcal{AS}}\left(\mathbf{W}^{0}_{\mathcal{SS}}-\mathbf{\Phi}\right)^{-1}\mathbf{W}^{0}_{\mathcal{SA}}\right]^{-1}.$ (31)
Using Eq. (30) and Eq. (29a), we obtain
$\left[\mathbf{\Omega}(\mathbf{c})\right]_{\mathcal{AA}}=\Big{[}\mathbf{\Omega}^{0}_{\mathcal{AA}}+\mathbf{W}^{0}_{\mathcal{AS}}\mathbf{\Sigma}\left(-\mathbf{\Phi}+\mathbf{\Sigma}\right)^{-1}\mathbf{\Sigma}\mathbf{W}^{0}_{\mathcal{SA}}\Big{]}^{-1}.$ (32)
Applying the Woodburry matrix identity to Eq. (32) and using Eq. (29b) and Eq. (29c) yields
$\left[\mathbf{\Omega}(\mathbf{c})\right]_{\mathcal{AA}}=\mathbf{\Omega}^{0}_{\mathcal{AA}}+\mathbf{\Omega}^{0}_{\mathcal{AS}}\left(\mathbf{\Phi}-\mathbf{\Omega}^{0}_{\mathcal{SS}}\right)^{-1}\mathbf{\Omega}^{0}_{\mathcal{SA}}.$ (33)
Finally, inserting Eq. (33) into Eq. (15) yields Eq. (16).
@@ Acknowledgment
The author acknowledges stimulating discussions with L. Le Magoarou, M. Nerini, H. Prod’homme, and A. Shaham.